OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2,8,-26,-80, … -2,8,-26,80, … -2,12,-44,180,… -2,-12,-44,-180, …

OpenStudy (freckles):

well the first two terms are already given -2,8,... we can find the third term by replacing n with 2

OpenStudy (freckles):

\[f(n)=-4f(n-1)-3f(n-2) \\ f(2)=-4f(2-1)-3f(2-2) \\ f(2)=-4f(1)-3f(0)\] now plug in the values you were given for f(0) and f(1) and apply order of operations to find f(2)

OpenStudy (anonymous):

i am lost... lol

OpenStudy (freckles):

do you know what f(1) and f(0) equal?

OpenStudy (anonymous):

8, -2

OpenStudy (freckles):

well plug those in then

OpenStudy (freckles):

\[f(n)=-4f(n-1)-3f(n-2) \\ f(2)=-4f(2-1)-3f(2-2) \\ f(2)=-4f(1)-3f(0) \\ f(2)=-4(8)-3(-2)\] do you know order of operations?

OpenStudy (anonymous):

not really, can you explain it to me?

OpenStudy (freckles):

1) exponents 2) grouping symbols 3) mult/divide as it occurs left to right 4) add/subtract as it occurs left to right

OpenStudy (freckles):

you only have multiplication and subtraction -4(8)=? 3(-2)=? then subtract the bottom product from the top

OpenStudy (anonymous):

-2 - (-32) -2 + 32 30

OpenStudy (freckles):

3(-2) is -6

OpenStudy (freckles):

-4(8)=-32

OpenStudy (freckles):

-32-(-6)=?

OpenStudy (anonymous):

-26

OpenStudy (freckles):

\[f(n)=-4f(n-1)-3f(n-2) \\ f(2)=-4f(2-1)-3f(2-2) \\ f(2)=-4f(1)-3f(0) \\ f(2)=-4(8)-3(-2) \\ f(2)=-32-(-6) \\ f(2)=-32+6 \\ f(2)=-26 \] ok use \[f(n)=-4f(n-1)-3f(n-2) \] again to find the 4th term in the sequence called f(3) f(3) so that means we want to replace n's with 3 now

OpenStudy (anonymous):

f(3)=−4f(3−1)−3f(3−2)

OpenStudy (freckles):

right so far

OpenStudy (freckles):

and 3-1=2 and 3-2=1 so you have f(3)=-4f(2)-3f(1)

OpenStudy (freckles):

you already know f(1) and you just found f(2)

OpenStudy (anonymous):

f(3)=-4(-26)-3(8)

OpenStudy (freckles):

yep now just apply the order of operations

OpenStudy (anonymous):

f(3)=-104-24?

OpenStudy (freckles):

almost keep in mind a neg times a neg=pos

OpenStudy (freckles):

so f(3)=104-24

OpenStudy (anonymous):

sorry my computer froze

OpenStudy (anonymous):

f(3)=80

OpenStudy (freckles):

great do you see your sequence

OpenStudy (anonymous):

yes i do :) thanks for all your help freckles! greatly appreciate it :)

OpenStudy (freckles):

you had f(0)=-2 f(1)=8 f(2)=-26 f(3)=80 so your sequence should look like this -2,8,-26,80,... and great work! :)